# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

### v2

@@ 8 Summary and Discussion
We introduce a new challenge of both scientific and practical value: decoding of arbitrary information seeking goals over specific texts from eye movements in reading. We tackle this challenge in two formulations, goal selection and goal reconstruction, using a number of evaluation frameworks and modeling approaches. The best performing model on the selection task is able to predict the correct question with a considerable degree of success, even when the candidate reading goals pertain to the same information in the text. Our results further suggest that although the reconstruction task is extremely challenging, meaningful progress can also be made in this open-ended regime. Overall, we find that despite the inherent noise in eye movement data, effective modeling can extract highly valuable information regarding specific information seeking goals at the challenging granularity level of eye movements of a single reader over a single paragraph.
Automatic decoding of information seeking goals in real time can pave the way for new applications with positive societal value. In the future, media outlets and providers of municipal and governmental services could better understand the information needs of users who access their websites and render information in a manner that better meets these needs. Automatic identification of information seeking goals can also be used for real-time assistance to special populations, like elderly people, when accessing critical information on the web. Future e-learning systems could assess students’ progress on various types of information-seeking tasks and monitor their information-seeking skills over time. Finally, one can imagine interactive reading interfaces that include real-time text personalization, such as simplification and suggestion of additional relevant information, according to the reader’s information seeking goals. The present study is a first step in enabling such applications. Ample room remains for further improvement of prediction accuracy via modeling innovations, as well as the collection and analysis of data for additional variants of information seeking regimes, corpora, reader populations, and languages.
@@ 9 Ethical Considerations
The OneStop Eye Movements dataset used in this work was collected by Berzak et al. (2025). The study was conducted under an institutional IRB protocol, and all the participants provided written consent before participation. The data is anonymized. Analyses and predictive modeling of task-based reading are among the primary use cases for which the data was collected.
@@ 10 Reproducibility Statement
We describe the model training and selection procedure, the evaluation protocol, the hyperparameter search space for each model, and the hardware and software specifications in Sections˜5 and C. The OneStop dataset (Berzak et al., 2025) used in the experiments and described in Section˜2 is publicly available at https://osf.io/2prdq/. The code for the paper, which implements all the models, experimental procedures, and analyses, is available at: https://anonymous.4open.science/r/open-question-prediction/.
@@
Appendix A Discriminative Models
@@ A.1 Haller RNN
    Fixation Sequence: Paragraph-only token-level embeddings are pooled to word-level embeddings, and ordered by the fixation sequence. Then, each embedding in the sequence is concatenated with the respective fixation features. The fixation-level features are:
@@
    • Total gaze duration (sum of all fixations on the word)
@@
Figure˜5(a) depicts the architecture adapted to our task.
A.2 RoBERTEye-Fixations
@@
    Fusion and Prediction: The combined sequence $Z_{E_{T}};\texttt{SEP}_{E};Z_{W}]$ is processed by the transformer encoder, and the final CLS token representation is passed through two fully connected layers to predict the question: $\hat{q}=\operatorname{argmax}\;\text{MLP}(\texttt{CLS})$.
Figure˜5(b) depicts the adapted architecture.
Table 3: Eye movement and word property features used by RoBERTEye-Fixations and DalEye-Llama. See Berzak et al. (2025) for further details.
@@
IA_SELECTIVE_REGRESSION_
PATH_DURATION
Duration of fixations and refixations on the current
interest area until eyes move to a higher ID.
@@
total_skip Binary indicator whether the word was fixated on.
Fixation-level Eye Movement Features
@@
Figure 5: Visualization of the Haller RNN and RoBERTEye-Fixations architectures. LMLM stands for a language model, RNNRNN for a recurrent neural network. $FF_{f_{i}}$ stands for the fixation features and $w_{f_{i}}$ for the word corresponding to the $i$-th fixation respectively.
Appendix B Generative Models
B.1 DalEye-LLaVA
DalEye-LLaVA extends the LLaVA-OneVision architecture by replacing the vision encoder with a fixation encoder tailored to eye movement features. The model input is structured as an instruction-style prompt comprising three components: (1) a task description, (2) the paragraph text, and (3) the participant’s eye movement features, positionally aligned with the paragraph words. The prompt concludes with the ground-truth question followed by an <eos> token. The entire sequence is tokenized, and training uses teacher forcing, with the cross-entropy loss applied only to the question tokens.
The forward architecture mirrors LLaVA’s design, with three main stages, as depicted in Figure˜6:
  1. 1.
    Text encoding: The task prompt and paragraph tokens are embedded by the backbone language model.
  2. 2.
    Fixation encoding: Eye movement features (see Table˜3) are processed by a dedicated fixation encoder, which applies MLP layers across the feature dimension and 1D convolutions across the temporal dimension, producing a sequence of fixation embeddings.
  3. 3.
B.2 DalEye-Llama
In Section˜B.2.1 we detail the task prompt and eye-movement input representations used for question generation with DalEye-Llama. Variants of these prompts and inputs are described in Section˜B.2.2.
B.2.1 Input format
@@
You will receive the eye movements data for the paragraph formatted as <FORMAT>
@@
<PARAGRAPH>
<SCANPATH>
<PARAGRAPH> is the textual paragraph, for example: ‘‘The quick brown fox jumps...’’.
<SCANPATH> is the eye movements data which depends on the <FORMAT>.
B.2.2 Input and Task Prompt Variations
@@
    Fixations: Fixation Duration + Next saccade direction
    • Format: a list of fixation-level features: [fixated word index, fixated word, fixation duration in ms, direction of next saccade (backward to earlier word / within word / forward to later word), words between current and next fixation]
    • Example: [[4, ‘‘fox’’, 220, backward], ...]
@@
    • Example: [[4, ‘‘fox’’, 320, 2, 1, 3, 0], ...]
@@
    Words+Fixations
We further experiment with an alternative prompt that de-emphasizes the experimental setup and places more focus on the model’s generation task:
Task Description:
Input Format:
You will receive the paragraph and eye movement data formatted as <FORMAT>
Expected Output:
Instructions:
Produce only the exact original question as your output.
<PARAGRAPH>
<SCANPATH>
Appendix C Model Training and Hyperparameters
@@ C.1 Discriminative Models
  • For RoBERTa-F, we search over dropout rates of $\{0.1,0.3,0.5\}$ for the eye movement projection layer. The model is trained with and without freezing language model parameters.
We train the deep-learning-based models for a maximum of 40 epochs, with early stopping after 8 epochs if no improvement in the validation error is observed. A single training epoch took roughly 5 minutes for RoBERTa-F, 10 minutes for Haller RNN, and 30 minutes for Llava and Llama. Each individual run was capped at 24 hours. Following Liu et al. (2019); Mosbach et al. (2021); Shubi et al. (2024), we use the AdamW optimizer (Loshchilov & Hutter, 2018) with a batch size of $16$. RoBERTa-F uses a linear warm-up ratio of $0.06$ and a weight decay of $0.1$. We standardize each eye movement feature using statistics computed on the training set, to zero mean unit variance.
@@
C.2 Generative Models
We present two generative models, DalEye-Llama and DalEye-LLaVA, each fine-tuned with distinct methods and hyperparameters optimized for their specific architectures and objectives.
@@ C.2 Generative Models
This model is fine-tuned using Unsloth with the Meta-Llama-3.1-8B backbone loaded in 4-bit precision. Training employs Low-Rank Adaptation (LoRA) with rank $r=16$, scaling factor $\alpha=16$, and applies RS-LoRA regularization. LoRA targets transformer modules: q_proj, k_proj, v_proj, up_proj, down_proj, o_proj, and gate_proj. The model is trained for 2 epochs with a linear scheduler and warm-up of 10 steps, batch size of 1, gradient accumulation over 2 steps, AdamW-8bit optimizer with learning rate $1\times 10^{-4}$, and weight decay of 0.01.
@@
This model employs a teacher-forcing strategy. Specifically, loss computation is restricted exclusively to tokens corresponding to possible questions, contrasting the likelihood of generating the correct question against two distractors. The computed losses for each question candidate are inverted, serving as logits for softmax normalization and subsequently optimized via cross-entropy against the true label. The LLaVA backbone remains frozen, and training employs LoRA with dropout rate of 0.1, RS-LoRA, and rank $r=8$. The hyperparameter search includes learning rates of $1\times 10^{-5},3\times 10^{-5},1\times 10^{-4}$}, fixation embedding hidden sizes of 512,1024{512,1024}, and early stopping after 8 epochs without validation improvement.
All neural network-based models were trained using the PyTorch Lightning (Falcon & team, 2024) library on NVIDIA A100-40GB and L40S-48GB GPUs.
We utilize the Hugging Face implementation of LLaVA-OneVision, specifically LLaVA-hf/LLaVA-onevision-qwen2-0.5b-si-hf. Additionally, we employ the gpt-4o-2024-08-06 version of GPT-4o. We use Unsloth (version 2025.9.7) (Daniel Han & team, 2023) to fine-tune the Llama model. For the QA model we use LIAMF-USP/roberta-large-finetuned-race.
There are roughly 355M trainable parameters for RoBERTa-F, reduced to 3M when the RoBERTa backbone is frozen. Haller RNN consists of 357M trainable parameters, or 2.4M when the backbone is frozen. DalEye-LLaVA has 505M parameters, out of which roughly 12M were unfrozen. DaLEye-Llama has 8.1B parameters, out of which roughly 42M were unfrozen.
Appendix D Question Reconstruction
D.1 Question Annotation into UIUC Categories
This section provides additional information regarding the question annotation into UIUC categories. Specifically, in Section˜D.1.1 we include the full prompt given to the model, in Section˜D.1.2 examples of questions and their corresponding annotations, and in Section˜D.1.3 agreement with human annotators.
D.1.1 Full Prompt
@@
D.1.2 Categorization Examples
For each UIUC question category we present an example question sourced from OneStopQA Berzak et al. (2020):

### v3

@@ 8 Summary and Discussion
We introduce a new challenge of both scientific and practical value: decoding arbitrary information seeking goals from eye movements in reading. We tackle this challenge in two formulations, goal selection and goal reconstruction, using a number of evaluation frameworks and modeling approaches. The best performing model on the selection task is able to predict the correct question with a considerable degree of success, even when the candidate reading goals pertain to the same information in the text. The performance of the model is driven by cognitively interpretable factors of human information seeking. Our results further suggest that although the reconstruction task is extremely challenging, meaningful progress can also be made in this open-ended regime. Overall, we find that effective modeling can extract highly valuable information regarding specific information seeking goals at the challenging granularity level of a single paragraph.
Automatic decoding of information seeking goals can pave the way for new applications with positive societal value. In the future, media outlets and providers of municipal and governmental services could better understand the information needs of users who access their websites and render information in a manner that better meets these needs. Automatic identification of information seeking goals can also be used for real-time assistance to special populations, when accessing critical information on the web. Future e-learning systems could assess students’ progress on information-seeking tasks and monitor their information-seeking skills over time. Finally, one can imagine interactive reading interfaces that include real-time text personalization, such as simplification and suggestion of additional relevant information, according to the reader’s information seeking goals. The present study is a first step in enabling such applications.
@@ 9 Ethical Considerations
The OneStop Eye Movements dataset used in this work was collected by Berzak et al. (2025b). The study was conducted under an institutional IRB protocol, and all the participants provided written consent before participation. The data is anonymized. Analyses and predictive modeling of task-based reading are among the primary use cases for which the data was collected.
@@ 11 Reproducibility Statement
We describe the model training and selection procedure, the evaluation protocol, the hyperparameter search space for each model, and the hardware and software specifications in Sections˜5 and D. The OneStop dataset (Berzak et al., 2025b) used in the experiments and described in Section˜2 is publicly available at https://osf.io/2prdq/. The code for the paper, which implements all the models, experimental procedures, and analyses, is available at: https://github.com/lacclab/Open-Ended-Goal-Decoding.
@@
Appendix A N-Gram Overlap Between Questions and Texts
To assess the extent to which the question answering task in OneStop (Berzak et al., 2025b) could be solved through simple lexical matching, we computed the n-gram overlap between each target question and its corresponding passage in OneStopQA (Berzak et al., 2020). We measured this overlap using the proportion of matching unigrams (ROUGE-1333We use ROUGE metrics (Lin, 2004), implemented by the rouge python package, version 1.0.1.), bigrams (ROUGE-2), and unigrams restricted to content words, with respect to the number of n-grams in the question (precision), in the text (recall), and the harmonic mean of the two ($F_{1}$). We further provide a breakdown of the question overlap with the passage into the words within and outside the question’s critical portion of the text. We compared this overlap with the corresponding overlap in SQuAD (Rajpurkar et al., 2016), an extractive QA dataset in which answers appear verbatim in the text. The results can be seen at table˜2.
Table 2: Average overlap results for OneStop and SQuAD, comparing questions to different parts of the context. Values in brackets refer to the analysis using only content words from both the question and the context. P (precision), R (recall), and $F_{1}$ refer to overlap proportion with respect to question length, context length, and their harmonic mean, respectively.
Dataset Text Part ROUGE-1 ROUGE-2
P R $F_{1}$ P R $F_{1}$
SQuAD Paragraph 0.527 (0.488) 0.071 (0.052) 0.122 (0.091) 0.208 0.021 0.037
OneStop Paragraph 0.463 (0.376) 0.059 (0.040) 0.104 (0.071) 0.127 0.012 0.022
CS 0.338 (0.285) 0.126 (0.097) 0.176 (0.137) 0.100 0.033 0.048
out of CS 0.283 (0.153) 0.050 (0.024) 0.084 (0.040) 0.034 0.005 0.008
We found that the unigram overlap between OneStop questions and passages (46%) is lower than the corresponding overlap in SQuAD (53%). We similarly find a smaller overlap in OneStop for content words (38% OneStop, 49% SQuAD), and for bigrams (12% OneStop, 20% SQuAD). We note that these comparisons are adequate with the ROUGE measures as the OneStop and SQuAD datasets have similar passage lengths on average (110 OneStop, 120 SQuAD) and question lengths (10 words on average in both). We further found that the n-gram overlap of the questions in OneStop with the critical span text is relatively low (33% for unigrams, 10% bigrams), and only slightly higher than with the text outside of the critical span (28% unigrams, 3% bigrams). This further suggests that the relevant text does not differ drastically from non-relevant text in terms of substring overlap with the question. Overall, we find that the overlap in OneStopQA is relatively low, both in absolute terms and compared to SQuAD, providing empirical support for the inferential nature of the information seeking tasks in OneStop.
Appendix B Discriminative Models
@@ B.1 Haller RNN
    Fixation Sequence: Paragraph-only token-level embeddings are sum-pooled to word-level embeddings, and ordered by the fixation sequence, including repetitions of words according to the number of fixations on that word. Then, each embedding in the sequence is concatenated with the respective fixation features. The fixation-level features are:
@@
    • Total gaze duration (sum of all fixations on the word; repeated for each fixation on that word)
@@
Figure˜6(a) depicts the architecture adapted to our task.
B.2 RoBERTEye-Fixations
@@
    Fusion and Prediction: The combined sequence $Z_{E_{T}};\texttt{SEP}_{E};Z_{W}]$ is processed by the transformer encoder, where $\texttt{SEP}_{E}$ is a separator token initialized the same as SEP. The final CLS token representation is passed through two fully connected layers to score each candidate question.
Figure˜6(b) depicts the adapted architecture.
Table 3: Eye movement and word property features used by RoBERTEye-Fixations and DalEye-LLaVA. See Berzak et al. (2025b) for further details.
@@
IA_SELECTIVE_REGRESSION_ PATH_DURATION Duration of fixations and refixations on the current interest area until eyes move to a higher ID.
@@
total_skip Binary indicator whether the word was fixated on.
Paragraph-level Features
@@
Fixation-level Fixation Features
@@
Figure 6: Visualization of the Haller RNN and RoBERTEye-Fixations architectures. LMLM stands for a language model, RNNRNN for a recurrent neural network. $FF_{f_{i}}$ stands for the fixation features and $w_{f_{i}}$ for the word corresponding to the $i$-th fixation.
Appendix C Generative Models
C.1 Prompts for DalEye-Llama, DalEye-GPT and Gemini
In Section˜C.1.1, we present the task prompt and eye-movement input representations used for question generation with DalEye-Llama, DalEye-GPT, and Gemini. Variants of these prompts and inputs are described in Section˜G.2.3. The few-shot examples selection method is described in Section˜C.1.2.
C.1.1 Input format
@@
You will receive the eye movements data for the paragraph formatted as {FORMAT}
@@
{PARAGRAPH}
{SCANPATH}
{PARAGRAPH} is the textual paragraph, for example: ‘‘The quick brown fox jumps...’’.
{SCANPATH} is the eye movements data which depends on the {FORMAT} depicted below.
C.1.2 Gemini Few-shot Input
In the few-shot setting, we prepend 10 sampled input–output examples, each pairing an eye movement trajectory with its corresponding paragraph and question. Examples are selected to match the evaluation regime while preventing information leakage: for New Text, the examples come from the same participant but different articles; for New Participant, from the same paragraph but different participants; and for New Participant & Text, from different participants and different articles. The examples precede the test instance in the prompt as shown below.
    {SYSTEM_MESSAGE}
    <Example>
        <INPUT>
            Paragraph:
            {PARAGRAPH_EXAMPLE}
            Fixations_data:
            {SCANPATH_EXAMPLE}
        </INPUT>
        <OUTPUT>
            {EXAMPLE_TRUE_QUESTION}
        </OUTPUT>
    </Example>
    ... # X 10 examples
        Paragraph:
        {PARAGRAPH_TEST}
        Fixations_data:
        {SCANPATH_TEST}
{SYSTEM_MESSAGE} is the same input prompt as in section C.1.1 (excluding the paragraph and scanpath data).
The four parameters with name format {PARAGRAPH/SCANPATH_EXAMPLE/TEST} are formatted as the {PARAGRAPH} and {SCANPATH} in section C.1.1 for each few-shot example and the test instance, correspondingly. {EXAMPLE_TRUE_QUESTION} is the corresponding question for each example.
C.2 DalEye-LLaVA
This model is based on the LLaVA-OneVision language-vision model architecture (Li et al., 2025). The model input consists of a task prompt, the text, and eye movement feature embeddings for the text words. We keep Qwen 2-0.5B (Yang et al., 2024) as the language model backbone, and replace the vision encoder with a novel trainable eye movement encoder, which encodes eye movement features using fully connected and convolutional layers. The eye movement embeddings are positionally aligned with the text words. The model is fine-tuned using an autoregressive next-token prediction objective for the correct question. Additional details, including a model diagram and eye movement features, are provided below. The model input is structured as an instruction-style prompt comprising three components: (1) a task description, (2) the paragraph text, and (3) the participant’s eye movement features, positionally aligned with the paragraph words. The prompt concludes with an <eos> token. The entire sequence is tokenized, and training uses teacher forcing, with cross-entropy loss applied to the ground-truth question tokens.
The forward architecture mirrors LLaVA’s design, with three main stages, as depicted in Figure˜7:
@@ C.2 DalEye-LLaVA
    Text encoding: The task prompt and paragraph are tokenized and embedded by the backbone language model.
@@
    Fixation encoding: Eye movement features (defined in Table˜3) are processed by a dedicated fixation encoder. The encoder first applies a two-layer MLP with LeakyReLU activation and dropout over the feature dimension (Linear $\rightarrow$ LeakyReLU $\rightarrow$ Dropout $\rightarrow$ Linear $\rightarrow$ LeakyReLU) and then applies two temporal 1D convolution layers (kernel size $3$, stride $1$, padding $1$), with LeakyReLU after each convolution and dropout between the two convolutions. This yields a sequence of fixation embeddings.
@@
Appendix D Model Training and Hyperparameters
@@ D.1 Discriminative Models
  • For RoBERTEye-Fixations, we search over dropout rates of $\{0.1,0.3,0.5\}$ for the eye movement projection layer. The model is trained with and without freezing language model parameters.
We train the deep-learning-based models for a maximum of 40 epochs, with early stopping after 8 epochs if no improvement in the validation error is observed. A single training epoch took roughly 5 minutes for RoBERTEye-Fixations, 10 minutes for Haller RNN, and 30 minutes for DalEye-Llama and DalEye-LLaVA. Each individual run was capped at 24 hours. Following Liu et al. (2019); Mosbach et al. (2021); Shubi et al. (2024), we use the AdamW optimizer (Loshchilov and Hutter, 2018) with a batch size of $16$. RoBERTEye-Fixations uses a linear warm-up ratio of $0.06$ and a weight decay of $0.1$. We standardize each eye movement feature using statistics computed on the training set, to zero mean unit variance.
@@
D.2 Generative Models
@@
This model is fine-tuned using Unsloth with the Meta-Llama-3.1-8B backbone loaded in 4-bit precision. Fine-tuning uses Low-Rank Adaptation (LoRA)(Hu et al., 2022) with all base model parameters frozen. We train only the LoRA adapters, configured with rank $r=16$, scaling factor $\alpha=16$, and RS-LoRA regularization. LoRA is applied to the transformer projection modules: q_proj, k_proj, v_proj, up_proj, down_proj, o_proj, and gate_proj. The model is fine-tuned for 2 epochs with a linear scheduler and warm-up of 10 steps, batch size of 1, gradient accumulation over 2 steps, AdamW-8bit optimizer with learning rate $1\times 10^{-4}$, and weight decay of 0.01.
DalEye-GPT
We fine-tune gpt-4o-mini-2024-07-18 using the OpenAI API for one epoch, with an automatically selected batch size of 4 and a learning-rate multiplier of 1.8.
@@
The LLaVA backbone remains frozen, and training employs LoRA with dropout rate of 0.1, RS-LoRA, and rank $r=16$. The hyperparameter search includes learning rates of $1\times 10^{-5},3\times 10^{-5},1\times 10^{-4}$}, fixation embedding hidden sizes of 512,1024{512,1024}, and early stopping after 7 epochs without validation improvement.
The discriminative models, DalEye-Llama, and DalEye-LLaVA were trained using the PyTorch Lightning (Falcon and team, 2024) library on NVIDIA A100-40GB and L40S-48GB GPUs.
We utilize the Hugging Face implementation of LLaVA-OneVision, specifically LLaVA-hf/LLaVA-onevision-qwen2-0.5b-si-hf. For GPT-4o we use version gpt-4o-2024-08-06. For Gemini 3 we use gemini-3-pro-preview version last updated on November 2025. We use Unsloth (version 2025.9.7) (Han et al., 2023) to fine-tune the Llama model. For the QA model we use LIAMF-USP/roberta-large-finetuned-race.
There are roughly 355M trainable parameters for RoBERTEye-Fixations, reduced to 3M when the RoBERTa backbone is frozen. Haller RNN consists of 357M trainable parameters, or 2.4M when the backbone is frozen. DalEye-LLaVA has 505M parameters, out of which roughly 12M were unfrozen. DalEye-Llama has 8.1B parameters, out of which roughly 42M were unfrozen.
Statistical testing is done with the Julia MixedModels package (Bates et al., 2022).
Appendix E Monty Hall-type Setup
Table˜4 illustrates the structure of the question selection task and the resulting chance accuracies under the three evaluation regimes. In the three-way All setting (a), the model selects among three candidate questions, yielding a chance accuracy of $3/9=33.3\%$. In the Different critical spans evaluation (b), success is relaxed to selecting either of the two questions associated with the critical span $c_{2}$ when it represents the correct span. Here, the probability of a correct random choice increases to $5/9=55.5\%$, analogous to a Monty Hall–type setup. Finally, in the Same critical span evaluation (c), the task reduces to a balanced binary choice between the two questions sharing $c_{2}$, resulting in a chance accuracy of $2/4=50\%$.
(a) All (b) Different critical spans (c) Same critical span
True $\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$}
$\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$}
$\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$}
Appendix F Question Reconstruction
F.1 Question Annotation into UIUC Categories
This section provides additional information regarding the question annotation into UIUC categories (Li and Roth, 2002). Specifically, in Section˜F.1.1 we include the full prompt given to the model, in Section˜F.1.2 examples of questions and their corresponding annotations, and in Section˜F.1.3 agreement with human annotators.
F.1.1 Full Prompt
@@
F.1.2 Categorization Examples
For each UIUC question category we present an example question sourced from OneStopQA (Berzak et al., 2020):